# Densities Of Primes And Primitive Roots

### v9

@@
Abstract: Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers.
@@ 1 Introduction
$\mathcal{P}_{u}=\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}\subset\mathbb{P}$ (1)
@@
$\pi_{u}(x)=\#\left\{p\leqslant x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (2)
@@
A conditional proof of this result was achieved in [23], and simplified sketches of the proof appear in [30, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [23, p. 218], [27], [28], et alii. An introduction to its historical development, and its calculations is covered in [30, pp. 3–10], and [42].
@@
$x^{-1}\sum_{u\leq x}\pi_{u}(x)=a_{0}\operatorname{li}(x)+O\left(x(\log x)^{-B}\right),$ (3)
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [18] unconditionally, and refined in [42]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (2) can be slightly different from the average quantity in (3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [23, p. 220], [30, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [40, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [38]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [19] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
$\mathcal{A}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 3\right\},$ (4)
@@
$\mathcal{B}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 1\right\},$ (5)
@@
$\pi_{u}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (6)
@@
Sections 2 to 6 introduce the notation and some standard results related to or applicable to the investigation of primitive roots in cyclic groups. The last section presents a proof of Theorem 1.1.
@@ 2.1 Simple Characters Sums
$\chi(u)=e^{i2\pi k\log u/(p-1)},$ (7)
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [29, p. 187], [34, p. 118], and [25, p. 271] for more details.
@@
$\psi(n)=e^{i2\pi kn/q},$ (8)
for some $k\geq 1,\gcd(k,q)=1$, see [29, p. 187], [34, p. 118], and [25, p. 271]. The additive character sums are quite similar to Lemma 2.1.
@@ 3.2 Divisors Dependent Characteristic Function
The works in [9], and [46] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [12, p. 863], [29, p. 258], [31, p. 18]. The characteristic function for multiple primitive roots is used in [7, p. 146] to study consecutive primitive roots. In [10] it is used to study the gap between primitive roots with respect to the Hamming metric. And in [46] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [9] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [38] using this formula.
@@ 3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [12], [10] for typical applications and [30, p. 19] for a discussion.
@@ 4 Estimates Of Exponential Sums
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p}+E(p),$ (14)
@@ 4.1 Incomplete And Complete Exponential Sums
Let $q=p+o(p)$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Applying the finite summation kernel in Definition 4.1, yields
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}=\sum_{n\leq x}\frac{1}{q}\sum_{0\leq t\leq q-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}.$ (20)
The term $t=0$ contributes x/q-x/q, and rearranging it yield
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}$ $=$ $\frac{1}{q}\sum_{n\leq x,}\sum_{1\leq t\leq q-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}-\frac{x}{q}$
$=$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right)\left(\sum_{n\leq x}\omega^{tn}\right)-\frac{x}{q}.$
@@
$\left|\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\right|$ $\leq$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left|\sum_{0\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{n\leq x}\omega^{tn}\right|+\frac{x}{q}$
$\ll$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left(2q^{1/2}\log q\right)\cdot\left(\frac{2q}{\pi t}\right)+\frac{x}{q}$
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\ll p^{1-\varepsilon}$ (24)
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}},$ (25)
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{1\leq n\leq p-1}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d).$ (26)
The term $t=0$ contributes $\varphi(p)/p$, and rearranging it yield
@@
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2qp{1/2}\log p\right)\cdot\left(\frac{4p^{1/2+\delta}}{\pi r(d,t)}\right)+\frac{\varphi(p)}{p}$
@@
where $r(d,t)\equiv dt\bmod p$ and $\delta>0$ is a small number. The last summation in (4.1) uses the estimate
$\sum_{1\leq t\leq p-1}\frac{1}{r(d,t)}=\sum_{1\leq z\leq p-1}\frac{1}{z}\ll\log p,$ (29)
since $z=r(d,t)\equiv dt\bmod p$ is a permutation in the finite field $\mathbb{F}_{p}$ and $\varphi(p)/p\leq 1$. This is restated in the simpler notation $p^{1/2+\delta}(\log p)^{2}\leq p^{1-\varepsilon}$ for any arbitrary small number $\varepsilon\in(0,1/2)$. ∎
The upper bound given in Theorem 4.2 seems to be optimum. A different proof, which has a weaker upper bound, appears in [15, Theorem 6], and related results are given in [6], [14], [20], and [21, Theorem 1].
@@ 4.2 Finite Fourier Transforms and Equivalent Exponential Sums
For any fixed $0\neq b\in\mathbb{F}_{p}$, the map $\tau^{n}\longrightarrow b\tau^{n}$ is one-to-one (permutation) in $\mathbb{F}_{p}$. Consequently, the subsets
$\{\tau^{n}:\gcd(n,p-1)=1\}\quad\text{ and }\quad\{b\tau^{n}:\gcd(n,p-1)=1\}\subset\mathbb{F}_{p}$ (30)
have the same cardinalities. As a direct consequence the exponential sums
$\sum_{\gcd(n,p-1)=1}e^{i2\pi b\tau^{n}/p}\quad\text{ and }\quad\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p},$ (31)
have the same upper bound up to an error term. An asymptotic relation for the finite Fourier transform of exponential sums (31) is provided here.
@@
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$ and $a\neq 0$ is not a primitive root, then
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+O(p^{1/2+\delta}(\log p)^{2}),$ (32)
@@
For $b\neq 1$, the exponential sum has the representation
$V(b)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p},$
confer equations (26) and (4.1) for more details. Similarly, for $b=1$,
$V(1)$ $=$ $\sum_{\gcd(n,p-1)=1}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p},$
@@
$V(b)-V(1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
@@
Taking the finite Fourier transform of the difference D(a)=V(b)V(1)D(a)=V(b)-V(1) returns
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
@@
By hypothesis $\tau^{s}-a\neq 0$ since $a\neq 0$ is not a primitive root. Thus, using the geometric sum identity $\sum_{1\leq u\leq p-1}e^{i2\pi u/p}=-1$ to simplify the last expression yields
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
@@
$\widehat{V(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ $=$ $\Bigg{|}\sum_{\gcd(n,p-1)=1}\omega^{tn}\Bigg{|}$ (39)
$\ll$ $\frac{p^{1+\delta}}{\pi r(d,t)},$
where $r(d,t)\equiv dt\bmod p$ and by Lemma 4.4, the difference including Gauss sum is bounded by
$\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}=\Bigg{|}1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg{|}$ $\leq$ $2p^{1/2}\log p,$ (40)
where $\chi(s)=e^{i\pi st/p}$, and $\psi(s)=e^{i2\pi\tau^{s}/p}$. Taking absolute value of the remainder term in (4.2) and replacing (39), and (40), return
$\widehat{R(a)}$ $=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}\cdot\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ (41)
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot\sum_{1\leq t\leq p-1}\frac{p^{1+\delta}}{\pi r(d,t)}$
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot\sum_{1\leq z\leq p-1}\frac{p^{1+\delta}}{\pi z}$

### v10

@@
Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. August 30, 2025
@@ 1 Introduction
$\mathcal{P}_{u}=\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}\subset\mathbb{P}$ (1.1)
@@
$\pi_{u}(x)=\#\left\{p\leqslant x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (1.2)
@@
A conditional proof of this result was achieved in [22], and simplified sketches of the proof appear in [29, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [22, p. 218], [26], [27], et alii. An introduction to its historical development, and its calculations is covered in [29, pp. 3–10], and [41].
@@
$x^{-1}\sum_{u\leq x}\pi_{u}(x)=a_{0}\operatorname{li}(x)+O\left(x(\log x)^{-B}\right),$ (1.3)
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [17] unconditionally, and refined in [41]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (1.2) can be slightly different from the average quantity in (1.3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [22, p. 220], [29, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [39, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [37]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [18] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
$\mathcal{A}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 3\right\},$ (1.4)
@@
$\mathcal{B}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 1\right\},$ (1.5)
@@
$\pi_{u}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (1.6)
@@
Section 2 to Section 6 introduce the notation and some standard results related to or applicable to the investigation of primitive roots in cyclic groups. The last section presents a proof of Theorem 1.1.
@@ 2.1 Simple Characters Sums
$\chi(u)=e^{i2\pi k\log u/(p-1)},$ (2.1)
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [28, p. 187], [33, p. 118], and [24, p. 271] for more details.
@@
$\psi(n)=e^{i2\pi kn/q},$ (2.2)
for some $k\geq 1,\gcd(k,q)=1$, see [28, p. 187], [33, p. 118], and [24, p. 271]. The additive character sums are quite similar to Lemma 2.1.
@@ 3.2 Divisors Dependent Characteristic Function
The works in [8], and [45] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [11, p. 863], [28, p. 258], [30, p. 18]. The characteristic function for multiple primitive roots is used in [6, p. 146] to study consecutive primitive roots. In [9], [25] it is used to study the gap between primitive roots with respect to the Hamming metric, et cetera. And in [45] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [8] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [37] using this formula.
@@ 3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [11], [9] for typical applications and [29, p. 19] for a discussion.
@@ 4 Estimates Of Exponential Sums
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p}+E(p),$ (4.1)
@@ 4.1 Incomplete And Complete Exponential Sums
Let $p$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Applying the finite summation kernel in Definition 4.1, yields
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}=\sum_{n\leq x}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}.$ (4.7)
Observe that the term $t=0$ contributes x/p-x/p, and rearranging it yield
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}$ $=$ $\frac{1}{p}\sum_{n\leq x,}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}-\frac{x}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq q-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right)\left(\sum_{n\leq x}\omega^{tn}\right)-\frac{x}{p}.$
@@
$\left|\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\right|$ $\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{0\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{n\leq x}\omega^{tn}\right|+\frac{x}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(p^{1/2}\log p\right)\cdot\left(\frac{2p}{\pi t}\right)+\frac{x}{q}$
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\ll p^{1-\varepsilon}$ (4.11)
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}},$ (4.12)
@@
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{1\leq n\leq p-1}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d).$ (4.13)
Now, observe that the term $t=0$ contributes $\varphi(p-1)/p$, and rearranging it yield
@@
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2p^{1/2}\log p\right)\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\sum_{1\leq t\leq p-1}\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
$\ll$ $\frac{1}{p}\left(2p^{1/2}\log p\right)\cdot\left(2p^{1+\delta}\log p\right)$
@@
where $\delta>0$ is a small number. This is restated in the simpler notation $p^{1/2+\delta}(\log p)^{2}\leq p^{1-\varepsilon}$ for any arbitrary small number $\varepsilon\in(0,1/2)$. ∎
The upper bound given in Theorem 4.2 seems to be optimum. A different proof, which has a weaker upper bound, appears in [14, Theorem 6], and related results are given in [5], [13], [19], [3], [2], [4], [14], [12] and [20, Theorem 1].
@@ 4.2 Finite Fourier Transforms and Equivalent Exponential Sums
An asymptotic relation for the finite Fourier transform of the exponential sum
$V(1,b)=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p},$ (4.16)
in terms of simpler exponential sum up to an error term, is provided here.
@@
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$, then
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+O(p^{1/2+\delta}(\log p)^{2}),$ (4.17)
@@
For $b\geq 1$, the exponential sum has the representation
$V(1,b)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p},$
confer equations (4.12), (4.13) and (4.1) for more details. In particular, for $b=1$,
$V(1,1)$ $=$ $\sum_{\gcd(n,p-1)=1}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p-1)}{p},$
@@
$V(1,b)-V(1,1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
@@
Let $a\in[1,p-1]$. Taking the finite Fourier transform of the difference D(a,b)=V(1,b)V(1,1)D(a,b)=V(1,b)-V(1,1) returns
$\widehat{D(a,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
@@
Now at $a=1$, $\tau^{s}-a\neq 0$ for any $s\in[1,p-1]$. Thus, using the geometric sum identity $\sum_{1\leq u\leq p-1}e^{i2\pi u/p}=-1$ to simplify the last expression yields
$\widehat{D(1,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
@@
$\widehat{V(1,b)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{b}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
@@
$\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ $=$ $\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{\gcd(n,p-1)=1}\omega^{tn}\Bigg{|}$ (4.24)
$\ll$ $p^{1+\delta}\log p,$
where $\delta>0$ is a small number and by Lemma 4.4, the difference including Gauss sum is bounded by
$\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}=\Bigg{|}1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg{|}$ $\leq$ $2p^{1/2}\log p,$ (4.25)
where $\chi(s)=e^{i\pi st/p}$, and $\psi(s)=e^{i2\pi\tau^{s}/p}$. Taking absolute value of the remainder term in (4.2) and replacing (4.24), and (4.25), return
[…]
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}\cdot\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot(p^{1+\delta}\log p)$